OpenStudy (virnios):

The dress store is having a sale where all merchandise is 1/4 off. A woman buys $48 of merchandise at the sale price. What would she have had to pay at the regular price.

OpenStudy (virnios):

Will medal, I really need help as soon as possible :D

OpenStudy (emmynimmy):

64 dollars

OpenStudy (virnios):

Thank you! Care to help me with 3 more questions? :P

OpenStudy (emmynimmy):

I can try

OpenStudy (virnios):

Mitch is buying candy bars for his friends. He wants to give 2 bars to each friend, and he wants to have 10 spare bars. He can afford to buy 24 candy bars. What is the algebraic sentence that will help him figure out how many friends he can terat? Let f = the number of friends. A. 24 + 10 = 24 B 1/2f - 10 = 24 C. 2f - 10 = 24 D. 1/2f + 10 = 24

OpenStudy (emmynimmy):

2f+10=24

OpenStudy (virnios):

Frank has devised a formula for his catering business that calculates the number of meatballs he needs to prepare. The formula is m = 4a + 2c, where c = number of children, m = number of meatballs, a = number of adults. How many meat balls are required for a party of 25 adults and 6 children?